OpenStudy (anonymous):

PLEASE HELP ME!! Given that EFGH is a kite, if FG = HG, find HG. A6 B18 C24 D not enough info.

OpenStudy (anonymous):



OpenStudy (anonymous):

@JA1 you want to help her on this one??

OpenStudy (anonymous):

OOPS f is across from H.

OpenStudy (anonymous):

Yes please!

OpenStudy (anonymous):

draw again please?

OpenStudy (anonymous):

i got a 72% on my last quiz thanks to you!! And sure thing.

OpenStudy (anonymous):

i only helped with one question.

OpenStudy (ja1):

Geometry? I pass my brain doesn't get along with this... too many formulas :3

OpenStudy (anonymous):



OpenStudy (anonymous):

Given that EFGH is a kite, if FG = HG, find HG. A6 B18 C24 D not enough info.

OpenStudy (anonymous):

still i probably would have failed wit hout that one question!!

OpenStudy (anonymous):

just writing the question

OpenStudy (anonymous):

so x+12 = x+12?? since their congruent to each other?

OpenStudy (anonymous):

I believe that the first 2 equations are 2x - 5 = x + 1 so that x = 6 then plug in the x into x + 12 = ??? you get the answer.

OpenStudy (anonymous):

i mean plug 6 into x + 12

OpenStudy (anonymous):

ohh! 7?

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

yes!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

The answers B!! YOU ARE A GENIUS!!

OpenStudy (anonymous):

thanks good luck!!

OpenStudy (anonymous):

thanks! i still have 50 minutes on my new quiz.!